OpenStudy (anonymous):

suppose that sec(t) =3/2 and that t is in 4th quadrant. find the exact value of tan(t)

OpenStudy (anonymous):



OpenStudy (anonymous):

that is (sqrt of 13)/2

OpenStudy (anonymous):

How did you arrive at this answer?

OpenStudy (lalaly):

\[\sec^2t-1=\tan^2t\] \[\tan^2t=(\frac{3}{2})^2-1\]

OpenStudy (lalaly):

\[tant=\frac{\sqrt5}{2}\]\[\tan^2t=\frac{5}{4}\]